OpenStudy (anonymous):

if y varies inversely as x, and y=77 when x=55 find x when y=55

OpenStudy (anonymous):

y = k/x 77 = k/55 k = 77* 55 when y = 55 55 = 77* 55/ x 55x = 77*55 x = 77